Week of 060619

Open Actions from last week:

Chair: James

SMOD:

Monday:

Log: LFC_DB_ERROR on 1131/lfc001. Twiki down

New Actions:

  • Notify developers of LFC errors - DONE (Bug in code on rename of symlink)
  • Castor1 DS move tomorrow - Jan

Discussion:

Tuesday:

Log: C2CMS cluster down - out of LSF licences

New Actions:

  • Jan/Olof - solve the C2CMS problem -take the 20 drainging nodes in SC4 cluster out in order to give licences to c2cms cluster (~15 needed). If not possible, take nodes out of batch to get licences.
Discussion:

Wednesday

Log:

Actions:

Discussion:

Thursday

Log:

Actions:

Discussion:

Friday

Log:

Actions:

Discussion:

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2006-06-20 - unknown
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback